# How Breakable Is Privacy: Probing and Resisting Model Inversion Attacks in Collaborative Inference

### v4

@@ Abstract.
Collaborative inference, model inversion attacks, deep learning security and privacy, information theory
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: ACM Conference on Computer and Communications Security; October 13–17, 2025; xx, xxisbn: 978-1-4503-XXXX-X/25/10ccs: Security and privacy Domain-specific security and privacy architectures
1. Introduction
@@
The rapid advancement of deep learning has led to increasingly complex neural networks that achieve remarkable performance in tasks like image recognition and object detection. However, deploying such models on edge devices remains challenging due to limited computational resources (Huckelberry et al., 2024). Applications like autonomous driving, facial recognition, IoT surveillance, and remote healthcare require real-time processing, yet the limited computational capacity of edge devices makes standalone deep learning inference impractical (Shlezinger and Bajić, 2022). Cloud computing provides substantial computational power but faces latency, bandwidth, and privacy concerns when transmitting raw data. Collaborative inference (CI) (Shlezinger et al., 2021) mitigates these issues by partitioning neural networks into an edge model $f_{\text{edge}}$ and a cloud model $f_{\text{cloud}}$, where $f_{\text{edge}}$ extracts feature representations for $f_{\text{cloud}}$ to inference. This method enhances efficiency while maintaining real-time performance and is widely applied in UAVs (Qu et al., 2024), IoT systems (Shlezinger and Bajić, 2022), and private cloud computing (PCC) (Engineering and Architecture, 2024). With the rise of 5G, IoT, and AI, CI is poised for broader adoption.
@@
However, recent studies have revealed significant privacy risks in CI, where input data on edge devices can be reconstructed by model inversion attacks (MIAs) (He et al., 2019; Yang et al., 2022; Yin et al., 2023), as shown in Figure 1. Unlike centralized deployments, where adversaries recover training data through model APIs (Yuan et al., 2023; Struppek et al., 2022), CI often relies on shallow neural networks at the edge, producing feature representations that retain significant mutual information (MI) with input data (Azizian and Bajić, 2024). This exposes raw inputs to adversaries, who can reconstruct sensitive data, such as facial images directly from transmitted features (Ding et al., 2024).
@@
MIAs in CI can be categorized into maximum likelihood estimati-on-based (MLE-based) MIAs (He et al., 2019, 2020) and generative model-based (Gen-based) MIAs (Yin et al., 2023; He et al., 2020). MLE-based attacks optimize input data to minimize the loss between the edge model’s output and transmitted features but require white-box access, limiting practicality. Gen-based MIAs, on the other hand, use generative models to approximate the inverse mapping of the edge model, reconstructing inputs from features. Their success depends on data distribution alignment and the model’s ability to learn the inverse function. These privacy risks have driven extensive research on attacks and defenses, but existing methods still face notable limitations.
Figure 1. Collaborative inference and the threat of model inversion attacks.
[…]
Collaborative inference and the threat of model inversion attacks.
@@
Firstly, existing research lacks a comprehensive analysis of the fundamental factors influencing MIA effectiveness in CI. Current attack methods focus on improving data reconstruction under specific assumptions, leveraging innovative techniques to enhance reconstruction fidelity and generalization. In recent years, defenses have increasingly recognized the role of mutual information in MIAs and have shifted toward designing optimal information bottleneck (IB) approaches (Wang et al., 2024; Peng et al., 2022) to counter MIAs. However, these methods remain a methodological extension based on existing findings rather than advancing a fundamental theoretical understanding of MIAs. Consequently, their improvements are inherently constrained by existing technical limitations. Identifying the key factors that control MIA effectiveness and developing a systematic criterion for assessing attack difficulty would significantly deepen our theoretical insights and offer broader theoretical guidance for designing more effective defense strategies.
@@
Secondly, the trade-off between usability, privacy, and deployability in edge models is an underexplored and challenging area in current research. Encryption-based defenses receive little attention in CI due to resource constraints, real-time demands (Mireshghallah et al., 2021), and vulnerability to side-channel attacks (Hu et al., 2024). Mainstream MIA defense strategies in CI, constrained by a limited understanding of MIA effectiveness factors, primarily focus on reducing the mutual information $I(x,z)$ between inputs $x$ and extracted features $z$ to counter MIAs (Wang et al., 2024; Peng et al., 2022; Azizian and Bajić, 2024). However, mutual information estimation is inherently challenging due to its reliance on accurate probability density estimation, large-scale training data, and the computational overhead associated with optimization techniques (Wang et al., 2024; Duan et al., 2023), making the usability-privacy trade-off inefficient and difficult to guarantee. Moreover, some approaches compromise usability (He et al., 2020), increase computational and storage burdens on edge devices (Wang et al., 2024), and complicate further fine-tuning and optimization. Consequently, such methods are often less favored in practical applications.
@@
Thus, there is a pressing need for a novel approach grounded in an in-depth analysis of attack theories to provide interpretable and effective privacy protection while maintaining the usability and deployability of edge models. This need gives rise to three challenges. Challenge 1 (Difficulty assessment): How to establish a verifiable criterion for assessing MIA difficulty in CI? The fundamental factors determining MIA effectiveness remain poorly understood, making it difficult to develop a reliable assessment criterion. While MI’s role in MIAs is recognized, its accurate estimation for validation remains challenging. Challenge 2 (Methodological design): How to effectively and interpretably remove redundant information from transmitted features while preserving edge model usability? Although IB theory provides a theoretically sound approach, its practical implementation is limited by the technical constraints of MI estimation. Current methods are computationally expensive and lack flexibility, making fine-tuning and further optimization difficult after initial deployment. Challenge 3 (Practical constraint): How to balance usability and privacy while ensuring practical deployability? Certain defense strategies (Azizian and Bajić, 2024) inevitably increase computational and storage burdens on edge devices. Furthermore, when applied to edge models designed for simple feature extraction, these methods struggle to counter advanced Gen-based MIAs effectively.
@@
To overcome the above challenges, we proposed a criterion $D_{mia}$ to evaluate MIA difficulty in CI and proposed a lightweight privacy-preserving strategy called Information Sifting Funnel (SiftFunnel), featuring 3 key aspects. (1) Criterion basis: SiftFunnel offers interpretable defense principles based on $D_{mia}$ We analyze MIAs using formulaic principles, information theory, and Fano’s inequality, and derive an evaluation formula involving 4 critical factors: mutual information $I(x,z)$, conditional entropy $H(x\mid z)$, feature entropy $H(z)$, and effective information mean $\delta(z)$ (the average number of non-zero elements in $z$). Their impact on MIA performance is validated via the Mutual Information Neural Estimator (MINE) (Belghazi et al., 2018). (2) Usability-privacy trade-off: SiftFunnel effectively reduces redundant information while preserving edge model usability. Aligning with $D_{mia}$ to strengthen privacy protection, we aim to reduce $H(z)$ and $\delta(z)$ while increasing $H(x\mid z)$. Specifically, we incorporate distance correlation, Pearson correlation constraints, and an $l_{1}$-norm constraint on $z$ to enforce these objectives. To maintain usability, we integrate label smoothing and KL-divergence loss into the training process. This optimization enables the edge model to minify redundant information while ensuring the extraction of task-relevant features. (3) Lightweight design: SiftFunnel balances usability and privacy while light-weighting the edge model. We adopt a funnel-shaped edge model to reduce computational and storage overhead. Attention blocks enhance feature selection, while an upsampling module on the cloud side compensates for lost details, preserving inference accuracy. This design, guided by $D_{mia}$, also improves privacy by limiting feature $\delta(z)$ exposure.
@@
  • We propose the first MIA difficulty criterion in CI, identifying $I(x,z)$, $H(x|z)$, $H(z)$ and $\delta(z)$ as factors impacting attack effectiveness, deepening insights into MIAs.
  • We propose SiftFunnel, a privacy-preserving scheme for CI that constrains redundant transmitted features through linear and nonlinear correlation, ensuring model usability with label smoothing and loss function regularization.
@@
  • We conduct systematic experiments validating the proposed criterion and demonstrating the effectiveness of SiftFunnel in balancing usability, privacy, and deployability.
2. Background and Related Work
2.1. Collaborative inference
Collaborative inference (Shlezinger et al., 2021) has emerged as an effective solution. This approach divides the inference process into two stages: one part is performed locally on the edge device (which may consist of a single device or a chain of interconnected devices), while the remaining computation is uploaded to the cloud. Formally, this can be expressed as
(1) $f(x)=f_{e}(x)\circ f_{c}(f_{e}(x))$
@@ 2.1. Collaborative inference
Collaborative inference typically begins with training $f$ on centralized datasets in the cloud, then partitioning and deploying the $f_{e}$ to the edge. Alternatively, approaches like Split Learning  (Zhu et al., 2023) allow edge devices to participate in model training locally. Regardless of the training approach, the inference process remains consistent. As depicted in Figure 1, the edge model computes an intermediate representation, which is transmitted to the cloud for completion of the inference process. The final output is either returned to the edge device or retained in the cloud for further analysis.
@@
However, existing research (He et al., 2019) highlights that edge models are susceptible to MIAs, where adversaries reconstruct input data from transmitted features. Notably, in edge-based inference, computations can be distributed across multiple devices in a chained architecture (He et al., 2019), expressed as $f=f_{e1}\circ f_{e2}...\circ f_{c}$. Studies (He et al., 2020; Ho et al., 2024) show that deeper neural networks closer to the cloud transmit more complex, decision-focused features, making them more resistant to reconstruction. In contrast, the initial edge device generates features retaining more input information, making it highly susceptible to MIAs. This paper focuses on the MIA threat to the initial edge device, which employs a shallow neural network while the remaining computation occurs in the cloud. This common scenario in real-world applications presents a significant security risk due to the increased likelihood of input data reconstruction, highlighting the need for effective countermeasures.
2.2. Model inversion attack
@@
Current MIA techniques in CI can be broadly categorized into MLE-based MIAs and Gen-based MIAs, with related research progress described as follows.
@@
MLE-based MIAs. MLE-based MIA was first introduced by Fredrikson et al. (Fredrikson et al., 2015), who reconstructed training data from a multilayer perceptron (MLP) by minimizing the difference between target and confidence scores using gradient descent. Subsequent research (He et al., 2019) incorporated Total Variation (TV) (Rudin et al., 1992) loss as a regularization term to improve optimization, with these advanced techniques referred to as rMLE-based MIA. In this paper, both are collectively referred to as MLE-based MIA. In CI scenarios, MLE-based methods are particularly effective for reconstructing data from shallow neural networks, as the richer feature representations and lower complexity provide adversaries with detailed gradient information that is easier to optimize compared to the prediction vectors of deeper networks.
@@
Gen-based MIAs. This approach was introduced by Yang et al. (Yang et al., 2019), who reconstructed input data in facial recognition systems by using generative deep learning models $g$ to approximate the inverse mapping function of the target model $f^{-1}$, such that $g\approx f^{-1}$. By feeding prediction vectors into $g$, they successfully generated facial input images. Building on this approach, He et al. (He et al., 2019) extended it to collaborative inference, demonstrating that input data from edge models is more vulnerable to reconstruction, thereby highlighting the risks. Unlike MLE-based methods, Gen-based MIA does not require white-box access to the edge model. Instead, it trains an attack model by calculating losses based on transmitted features and auxiliary data, enabling it to approximate the inverse mapping of non-linear computations. Furthermore, as edge models produce rich and redundant outputs, this facilitates the training of attack models and the reconstruction process, as analyzed in detail in Section 4.
2.3. Related defense work
@@
To safeguard input data collected by edge devices in CI from potential leakage, He et al. (He et al., 2019, 2020) emphasized the importance of developing effective defense strategies while identifying these vulnerabilities. Existing defenses can be classified into three main categories: perturbation-based, IB-based, and neural network depth (NND) & IB-based approaches. The following sections outline these strategies.
@@
Perturbation-based defense. This approach introduces carefully designed noise perturbations into the gradients for training or trained edge model output features to reduce the likelihood of adversaries extracting meaningful information and defending against MIA. He et al. (He et al., 2020) initially proposed using Gaussian noise and random feature dropout to limit the effective information accessible to attackers, and later, Wang et al. (Wang et al., 2022) extended this strategy by dynamically adjusting the depth of the edge model while adding differential privacy noise to transmitted features. Although these methods enhance defenses, they significantly impact model usability, typically reducing accuracy by approximately 10%, and remain less effective against Gen-based MIA. The root cause is that the decoder in Gen-based MIAs is sufficiently expressive to adapt to noise or feature dropout, similar to the training of a denoising auto-encoder (DAE) (Gondara, 2016), potentially increasing attack robustness. In centralized MIA defenses, Struppek et al. (Struppek et al., 2023) proposed perturbing training labels with a negative label smoothing factor, significantly reducing MIA effectiveness and offering a novel perspective for designing defense strategies.
IB-based defense. This approach enhances defenses against inference attacks by limiting the mutual information between input data $x$ and intermediate features $z$ while preserving the mutual information between $z$ and outputs $y$, thus maintaining inference performance (Wang et al., 2021). The objective of this defense can be formally expressed as follows:
(2) $\min_{\theta}-{I}(z,y)+{I}(x,z)$
@@
Initially, Wang et al. (Wang et al., 2021) introduced information bottleneck theory into MIA defenses, but their method was not optimized for CI and required substantial computational resources for intermediate feature processing. To address this, Peng et al. (Peng et al., 2022) proposed BiDO (Bilateral Dependency Optimization), which uses dependency algorithms to optimize correlations among inputs, features, and outputs. BiDO reduces input-feature correlation while enhancing feature-output correlation, achieving a balance between defense and performance. Additionally, Wang et al. (Wang et al., 2024) and Duan et al. (Duan et al., 2023) explored mutual information estimation between $x$ and $z$ using methods based on mutual information definition and the CLUB (Contrastive Log Upper Bound) (Cheng et al., 2020) technique, incorporating the estimates as constraints in the training loss function. IB-based defenses provide interpretable protection mechanisms and valuable design insights for future research. However, these methods face limitations: BiDO relies on kernel functions for non-linear correlation constraints, with performance highly sensitive to hyper-parameters like kernel bandwidth; poor parameter selection can lead to suboptimal results. Methods based on mutual information estimation are constrained by sample scale to compute.
@@
NND & IB-based defense. Leveraging deeper networks and IB theory, this approach enhances defense effectiveness with targeted loss functions. Ding et al. (Ding et al., 2024) proposed increasing network depth while compressing newly added hidden layers to restrict redundant information transmission. Subsequently, Azizian et al. (Azizian and Bajić, 2024) further improved defenses by introducing an auto-encoder (AE) structure, with encoders at the edge and decoders in the cloud. The edge encoder increases network depth and reduces feature dimensionality to achieve the IB effect, while the cloud decoder restores information to maintain predictive performance while applying an $l_{1}$-norm constraint to reduce attack risks. Adversarial training (Ding et al., 2024) is also incorporated to strengthen robustness against MIAs.
@@
However, this method has certain limitations. Increasing network depth and adding modules raise computational and hardware demands on edge devices, limiting applicability. Adversarial training, while strengthening defenses, introduces additional challenges—it is less effective against advanced attacks, complicates fine-tuning, and degrades inference performance (Zhang et al., 2019). Additionally, adversaries can adapt their attack models using similar adversarial strategies, weakening defense effectiveness (Jin et al., 2024). While these methods improve defenses, their implementation requires careful consideration of computational resources, model adaptability, and application suitability. SiftFunnel enhances computational efficiency by deepening edge model structures and adopting a lightweight design, exhibiting IB-like characteristics after training. Thus, our method falls under the NND & IB-based defense strategy.
3. Threat Model
3.1. Knowledge background of adversary

### v5

@@ Abstract
1 Introduction
@@
The rapid advancement of deep learning has led to increasingly complex neural networks that achieve remarkable performance in tasks like image recognition and object detection. However, deploying such models on edge devices remains challenging due to limited computational resources [1]. Applications like autonomous driving, facial recognition, IoT surveillance, and remote healthcare require real-time processing, yet the limited computational capacity of edge devices makes standalone deep learning inference impractical [2]. Cloud computing provides substantial computational power but faces latency, bandwidth, and privacy concerns when transmitting raw data. Collaborative inference (CI) [3] mitigates these issues by partitioning neural networks into an edge model $f_{\text{edge}}$ and a cloud model $f_{\text{cloud}}$, where $f_{\text{edge}}$ extracts feature representations for $f_{\text{cloud}}$ to inference. This method enhances efficiency while maintaining real-time performance and is widely applied in UAVs [4], IoT systems [2], and private cloud computing (PCC) [5]. With the rise of 5G, IoT, and AI, CI is poised for broader adoption.
@@
However, recent studies have revealed significant privacy risks in CI, where input data on edge devices can be reconstructed by model inversion attacks (MIAs) [6, 7, 8], as shown in Figure 1. Unlike centralized deployments, where adversaries recover training data through model APIs [41, 53], CI often relies on shallow neural networks at the edge, producing feature representations that retain significant mutual information (MI) with input data [15]. This exposes raw inputs to adversaries, who can reconstruct sensitive data, such as facial images directly from transmitted features [20].
@@
MIAs in CI can be categorized into maximum likelihood estimati-on-based (MLE-based) MIAs [6, 16] and generative model-based (Gen-based) MIAs [8, 16]. MLE-based attacks optimize input data to minimize the loss between the edge model’s output and transmitted features but require white-box access, limiting practicality. Gen-based MIAs, on the other hand, use generative models to approximate the inverse mapping of the edge model, reconstructing inputs from features. Their success depends on data distribution alignment and the model’s ability to learn the inverse function. These privacy risks have driven extensive research on attacks and defenses, but existing methods still face notable limitations.
Figure 1: Collaborative inference and the threat of model inversion attacks.
@@
Firstly, existing research lacks a comprehensive analysis of the fundamental factors influencing MIA effectiveness in CI. Current attack methods focus on improving data reconstruction under specific assumptions, leveraging innovative techniques to enhance reconstruction fidelity and generalization. In recent years, defenses have increasingly recognized the role of mutual information in MIAs and have shifted toward designing optimal information bottleneck (IB) approaches [19, 18] to counter MIAs. However, these methods remain a methodological extension based on existing findings rather than advancing a fundamental theoretical understanding of MIAs. Consequently, their improvements are inherently constrained by existing technical limitations. Identifying the key factors that control MIA effectiveness and developing a systematic criterion for assessing attack difficulty would significantly deepen our theoretical insights and offer broader theoretical guidance for designing more effective defense strategies.
@@
Secondly, the trade-off between usability, privacy, and deployability in edge models is an underexplored and challenging area in current research. Encryption-based defenses receive little attention in CI due to resource constraints, real-time demands [22], and vulnerability to side-channel attacks [23]. Mainstream MIA defense strategies in CI, constrained by a limited understanding of MIA effectiveness factors, primarily focus on reducing the mutual information $I(x,z)$ between inputs $x$ and extracted features $z$ to counter MIAs [19, 18, 15]. However, mutual information estimation is inherently challenging due to its reliance on accurate probability density estimation, large-scale training data, and the computational overhead associated with optimization techniques [19, 34], making the usability-privacy trade-off inefficient and difficult to guarantee. Moreover, some approaches compromise usability [16], increase computational and storage burdens on edge devices [19], and complicate further fine-tuning and optimization. Consequently, such methods are often less favored in practical applications.
@@
Thus, there is a pressing need for a novel approach grounded in an in-depth analysis of attack theories to provide interpretable and effective privacy protection while maintaining the usability and deployability of edge models. This need gives rise to three challenges. Challenge 1 (Difficulty assessment): How to establish a verifiable criterion for assessing MIA difficulty in CI? The fundamental factors determining MIA effectiveness remain poorly understood, making it difficult to develop a reliable assessment criterion. While MI’s role in MIAs is recognized, its accurate estimation for validation remains challenging. Challenge 2 (Methodological design): How to effectively and interpretably remove redundant information from transmitted features while preserving edge model usability? Although IB theory provides a theoretically sound approach, its practical implementation is limited by the technical constraints of MI estimation. Current methods are computationally expensive and lack flexibility, making fine-tuning and further optimization difficult after initial deployment. Challenge 3 (Practical constraint): How to balance usability and privacy while ensuring practical deployability? Certain defense strategies [15] inevitably increase computational and storage burdens on edge devices. Furthermore, when applied to edge models designed for simple feature extraction, these methods struggle to counter advanced Gen-based MIAs effectively.
@@
To overcome the above challenges, we proposed a criterion $D_{mia}$ to evaluate MIA difficulty in CI and proposed a lightweight privacy-preserving strategy called Information Sifting Funnel (SiftFunnel), featuring 3 key aspects. (1) Criterion basis: SiftFunnel offers interpretable defense principles based on $D_{mia}$ We analyze MIAs using formulaic principles, information theory, and Fano’s inequality, and derive an evaluation formula involving 4 critical factors: mutual information $I(x,z)$, conditional entropy $H(x\mid z)$, feature entropy $H(z)$, and effective information mean $\delta(z)$ (the average number of non-zero elements in $z$). Their impact on MIA performance is validated via the Mutual Information Neural Estimator (MINE) [24]. (2) Usability-privacy trade-off: SiftFunnel effectively reduces redundant information while preserving edge model usability. Aligning with $D_{mia}$ to strengthen privacy protection, we aim to reduce $H(z)$ and $\delta(z)$ while increasing $H(x\mid z)$. Specifically, we incorporate distance correlation, Pearson correlation constraints, and an $l_{1}$-norm constraint on $z$ to enforce these objectives. To maintain usability, we integrate label smoothing and KL-divergence loss into the training process. This optimization enables the edge model to minify redundant information while ensuring the extraction of task-relevant features. (3) Lightweight design: SiftFunnel balances usability and privacy while light-weighting the edge model. We adopt a funnel-shaped edge model to reduce computational and storage overhead. Attention blocks enhance feature selection, while an upsampling module on the cloud side compensates for lost details, preserving inference accuracy. This design, guided by $D_{mia}$, also improves privacy by limiting feature $\delta(z)$ exposure.
@@
  • We propose the first MIA difficulty criterion in CI, identifying $I(x,z)$, $H(x|z)$, $H(z)$ and $\delta(z)$ as factors impacting attack effectiveness, deepening insights.
  • We propose SiftFunnel, a privacy-preserving scheme for CI that constrains redundant features via linear and nonlinear correlation, ensuring model usability with label smoothing and loss function regularization.
@@
  • We conduct systematic experiments to validate the proposed criterion and demonstrate SiftFunnel’s effectiveness in balancing usability, privacy, and deployability, using additional metrics including mutual information, $\delta(z)$, and parameter count for evaluation.
2 Background and Related Work
2.1 Collaborative inference
Collaborative inference [3] has emerged as an effective solution. This approach divides the inference process into two stages: one part is performed locally on the edge device (which may consist of a single device or a chain of interconnected devices), while the remaining computation is uploaded to the cloud. Formally, this can be expressed as
$f(x)=f_{e}(x)\circ f_{c}(f_{e}(x))$ (1)
@@ 2.1 Collaborative inference
CI typically begins with training $f$ on centralized datasets in the cloud, then partitioning and deploying the $f_{e}$ to the edge. Therefore, it is imperative to note that CI represents a scenario that has advanced to the inference application phase. Alternatively, approaches like Split Learning  [28] allow edge devices to participate in model training locally. Regardless of the training approach, the inference process remains consistent. As depicted in Figure 1, the edge model computes an intermediate representation, which is transmitted to the cloud for completion of the inference process. The final output is either returned to the edge device or retained in the cloud for further analysis.
@@
However, existing research [6] highlights that edge models are susceptible to MIAs, where adversaries reconstruct input data from transmitted features. Notably, in edge-based inference, computations can be distributed across multiple devices in a chained architecture [6], expressed as $f=f_{e1}\circ f_{e2}...\circ f_{c}$. Studies [16, 49] show that deeper neural networks closer to the cloud transmit more complex, decision-focused features, making them more resistant to reconstruction. In contrast, the initial edge device generates features retaining more input information, making it highly susceptible to MIAs. This paper focuses on the MIA threat to the initial edge device, which employs a shallow neural network while the remaining computation occurs in the cloud. This common scenario in real-world applications presents a significant security risk due to the increased likelihood of input data reconstruction, highlighting the need for effective countermeasures.
2.2 Model inversion attack
@@
Current MIAs in CI can be broadly categorized into MLE-based MIAs and Gen-based MIAs, with related research progress described as follows.
@@
MLE-based MIAs. MLE-based MIA was first introduced by Fredrikson et al. [29], who reconstructed training data from a multilayer perceptron (MLP) by minimizing the difference between target and confidence scores using gradient descent. Subsequent research [6] incorporated Total Variation (TV) [30] loss as a regularization term to improve optimization, with these advanced techniques referred to as rMLE-based MIA. In this paper, both are collectively referred to as MLE-based MIA. In CI scenarios, MLE-based methods are particularly effective for reconstructing data from shallow neural networks, as the richer feature representations and lower complexity provide adversaries with detailed gradient information that is easier to optimize compared to the prediction vectors of deeper networks.
@@
Gen-based MIAs. This approach was introduced by Yang et al. [10], who reconstructed input data in facial recognition systems by using generative deep learning models $g$ to approximate the inverse mapping function of the target model $f^{-1}$, such that $g\approx f^{-1}$. By feeding prediction vectors into $g$, they successfully generated facial input images. Building on this approach, He et al. [6] extended it to collaborative inference, demonstrating that input data from edge models is more vulnerable to reconstruction, thereby highlighting the risks. Unlike MLE-based methods, Gen-based MIA does not require white-box access to the edge model. Instead, it trains an attack model by calculating losses based on transmitted features and auxiliary data, enabling it to approximate the inverse mapping of non-linear computations. Furthermore, as edge models produce rich and redundant outputs, this facilitates the training of attack models and the reconstruction process, as analyzed in detail in Section 4.
Facts and Motivation. Although MIA may be confused with “Membership Inference Attack” [9], its meaning is clearly defined in this work and widely adopted in the model inversion literature. Moreover, the feasibility of MIA and their threat to CI have been well established. While encryption may protect data in transit, it does not prevent input reconstruction once intermediate features are exposed. Such exposure can occur locally or via compromised cloud infrastructure. Moreover, encryption adds substantial overhead to edge systems, limiting its practicality. These concerns highlight the need to analyze attack factors and design lightweight, deployable defenses.
2.3 Related defense work
@@
To safeguard input data collected by edge devices in CI from potential leakage, He et al. [6, 16] emphasized the importance of developing effective defense strategies while identifying these vulnerabilities. Existing defenses can be classified into three main categories: perturbation-based, IB-based, and neural network depth (NND) & IB-based approaches. The following sections outline these strategies.
@@
Perturbation-based defense. This approach introduces carefully designed noise perturbations into the gradients for training or trained edge model output features to reduce the likelihood of adversaries extracting meaningful information and defending against MIA. He et al. [16] initially proposed using Gaussian noise and random feature dropout to limit the effective information accessible to attackers, and later, Wang et al. [17] extended this strategy by dynamically adjusting the depth of the edge model while adding differential privacy noise to transmitted features. Although these methods enhance defenses, they significantly impact model usability, typically reducing accuracy by approximately 10%, and remain less effective against Gen-based MIA. The root cause is that the decoder in Gen-based MIAs is sufficiently expressive to adapt to noise or feature dropout, similar to the training of a denoising auto-encoder (DAE) [32], potentially increasing attack robustness. In centralized MIA defenses, Struppek et al. [33] proposed perturbing training labels with a negative label smoothing factor, significantly reducing MIA effectiveness and offering a novel perspective for designing defense strategies.
@@
IB-based defense. This approach enhances defenses against inference attacks by limiting the mutual information between input data $x$ and intermediate features $z$ while preserving the mutual information between $z$ and outputs $y$, thus maintaining inference performance [21]. The objective of this defense can be formally expressed as $\min_{\theta}-{I}(z,y)+{I}(x,z)$. Initially, Wang et al. [21] introduced information bottleneck theory into MIA defenses, but their method was not optimized for CI and required substantial computational resources for intermediate feature processing. To address this, Peng et al. [18] proposed BiDO (Bilateral Dependency Optimization), which uses dependency algorithms to optimize correlations among inputs, features, and outputs. BiDO reduces input-feature correlation while enhancing feature-output correlation, achieving a balance between defense and performance. Additionally, Wang et al. [19] and Duan et al. [34] explored mutual information estimation between $x$ and $z$ using methods based on mutual information definition and the CLUB (Contrastive Log Upper Bound) [35] technique, incorporating the estimates as constraints in the training loss function. IB-based defenses provide interpretable protection mechanisms and valuable design insights for future research. However, these methods face limitations: BiDO relies on kernel functions for non-linear correlation constraints, with performance highly sensitive to hyper-parameters like kernel bandwidth; poor parameter selection can lead to suboptimal results. Methods based on mutual information estimation are constrained by sample scale to compute.
@@
NND & IB-based defense. Leveraging deeper networks and IB theory, this approach enhances defense effectiveness with targeted loss functions. Ding et al. [20] proposed increasing network depth while compressing newly added hidden layers to restrict redundant information transmission. Subsequently, Azizian et al. [15] further improved defenses by introducing an auto-encoder (AE) structure, with encoders at the edge and decoders in the cloud. The edge encoder increases network depth and reduces feature dimensionality to achieve the IB effect, while the cloud decoder restores information to maintain predictive performance while applying an $l_{1}$-norm constraint to reduce attack risks. Adversarial training [20] is also incorporated to strengthen robustness against MIAs.
@@
However, this method has certain limitations. Increasing network depth and adding modules raise computational and hardware demands on edge devices, limiting applicability. Adversarial training, while strengthening defenses, introduces additional challenges—it is less effective against advanced attacks, complicates fine-tuning, and degrades inference performance [36]. Additionally, adversaries can adapt their attack models using similar adversarial strategies, weakening defense effectiveness [37]. While these methods improve defenses, their implementation requires careful consideration of computational resources, model adaptability, and application suitability. SiftFunnel enhances computational efficiency by deepening edge model structures and adopting a lightweight design, exhibiting IB-like characteristics after training. Thus, our method falls under the NND & IB-based defense strategy.
3 Threat Model
3.1 Knowledge background of adversary